40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 01:45:41 2011, Keiko, Update, LSC, Locking status update PRCL-OLT.pdfMICH-OLT.pdf
    Reply  Thu Sep 22 04:49:14 2011, Anamaria, Update, LSC, Locking status update - Some Scripts, No Louck 
    Reply  Thu Sep 22 11:50:37 2011, Koji, Update, LSC, Locking status update 
Message ID: 5516     Entry time: Thu Sep 22 11:50:37 2011     In reply to: 5512
Author: Koji 
Type: Update 
Category: LSC 
Subject: Locking status update 

Both loops basically have no phase margins. i.e. unstable. How can you lock PRMI with these servos?

Quote:

The following rough swept sine plots are the OLTs for MICH and PRCL. The gain setting was -10 and 0.5 for MICH and PRCL, respectively. Integrators were off. Looking at the measured plots, MICH has about 300 Hz UGF, when the gain is -20, and PRCL has about 300 HZ UGF, too, when the gain is 0.8.

 

ELOG V3.1.3-